Solve and check the linear equation.
3(x-6)-6(x-3)=x+3-(x-3)
The solution se is { }?

Answer by chen.aavaz(62)

3(x-6)-6(x-3)=x+3-(x-3)
3x-18-6x+18=x+3-x+3
-3x=6
x=-6/3
x=-2